OpenStudy (anonymous):

most insurance companies will replace a vehical anytime an estimate repaire exceeds 80 % of the "blue book" value if the vehical. Michelles insurance company paid $7500 for repaires on her car after an accident. what can be concluded about the blue book value of the car?

OpenStudy (mathstudent55):

This means that $7500 is at most 80% of the value of her vehicle. Do a proportion to find the full blue book value: Full value: 100% $7500: 80% v/100 = 7500/80 Solve for v (value): 80v = 100 x 7500 80v = 750,000 v = 9,375 That means her vehicle is worth at least $9,375